If (x-4) (x+3)=0 which is the greater of the two roots?

Answer by pahuja(2)

The quadratic equation in question is:
(x-4) (x+3)=0

Solution:
The roots of this quadratic are:
4 (x - 4 = 0, or x = 4)
and
-3 (x + 3 = 0, or x = -3)
The greater of the two roots is 4 ( 4 > -3)